OpenStudy (anonymous):

Does sin 2θ mean you are supposed to multiply the angle before calculating sin?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank yoooouuu

OpenStudy (anonymous):

ywwwww

OpenStudy (anonymous):

if you're still there, i think satellite may be wrong...

OpenStudy (anonymous):

it's true, you can multiply the angle, but i'm guessing your teacher wants you to use the trig identity \[\sin(2\theta)=2*\sin \theta * \cos \theta\]

OpenStudy (anonymous):

the next question says tan θ/2

OpenStudy (anonymous):

would that make sense for an identity as well?

OpenStudy (anonymous):

yup.

OpenStudy (anonymous):

hold on, i'll type the identity

OpenStudy (anonymous):

I have got a range that θ is in 90 degrees < θ < 180 degrees then it says find the sign for the following....

OpenStudy (anonymous):

sin 2θ

OpenStudy (anonymous):

\[\frac{ \sin \theta }{ \cos \theta + 1 }\] omg i suck at typing.

OpenStudy (anonymous):

the above's for the tan theta/2

OpenStudy (anonymous):

so I thought that meant 182<2θ<358

OpenStudy (anonymous):

OK thanks man I will see what I can work out.

OpenStudy (anonymous):

i think it would be negative for that

OpenStudy (anonymous):

Ok I think you are right there is a section about this in the notes he posted online.